OpenStudy (anonymous):

A coin is flipped 8 times. Find the probability of the event: exactly 3 heads

OpenStudy (anonymous):

@mebs

OpenStudy (anonymous):

So your in data management right. Do you know any standard distribution?... here's how you do these questions.

hero (hero):

(8 choose 3)(1/2)^3(1/2)^5

OpenStudy (anonymous):

the probability of heads is 1/2 and so is tails so \[8C3(\frac{ 1 }{ 2 })^{3} (\frac{ 1 }{ 2 })^{5}\] the three represents the heads the other five must be the tails.

OpenStudy (akashdeepdeb):

That would mean that if it rtossed thrice their can be 8 out comes (2^3) T T T H H H T T H H T H T H H H H T T H T H T T The probability of exactly 3 heads is = 1 possible out come/ 8 total outcomes = 1/8 Understood?

OpenStudy (anonymous):

dude you missed so many outcomes....

OpenStudy (anonymous):

the reason we just distribution patterns is not to do that kind of work.. though it is good to see it visually

hero (hero):

Actually, I agree with @mebs @AkashdeepDeb assumes the person flipped three coins 8 times, which is not the scenario the question is referring to.